OpenStudy (anonymous):

Evaluate the definite integral: ∫[0,1,x(x^(2)+1)^(5),dx] Please help!!! thanks!;)

OpenStudy (zehanz):

Is this it: \[\int\limits_{0}^{1}x(x^2+1)^5dx\] In that case, let u = x^2+1, then du = 2xdx, so the integral becomes:\[\frac{ 1 }{ 2 } \int\limits_{1}^{2}u^5du\] Can you evaluate this?

OpenStudy (anonymous):

Ok, Yep!! let me try...

OpenStudy (anonymous):

Nop, I got stcuk!!! urgh!! lol!

OpenStudy (anonymous):

Thanks I got it now.....It was hard though,...

OpenStudy (zehanz):

YW!